I have to solve this. I have to a point. Sharon has 30 nickels and dimes having a total value of $2.35. How many of each kind does she have? Fill in the chart the solve:
# of coins X Value per coin = Total Value
Nickels X 5 i
Dimes ii 10 iii
Total 30 2.35
i is 5x, ii would be 30-x but not sure what iii would be I think 10x maybe. I'm stuck on how to set up the equation then solve. Please help. thank you.

Answer by richwmiller(9135)

This is a word problem with coins.
What do we know ?
There are 30 coins and the coins are nickels and dimes
so n+d=30
we also know their value.
5n+10d=235
now we have two equations with two unknowns so we can solve it.
n+d=30
5n+10d=235
n=30-d
substitute
5(30-d)+10d=235
150-5d+10d=235
150+5d=235
5d=85
d=85/5
d=17
so n=13
check
5*13+17*10=235
65+170=235
235=235